# Towards Balanced Multi-Modal Learning in 3D Human Pose Estimation

### v4

@@ Abstract
3D human pose estimation (3D HPE) has emerged as a prominent research topic, particularly in the realm of RGB-based methods. However, the use of RGB images is often limited by issues such as occlusion and privacy constraints. Consequently, multi-modal sensing, which leverages non-intrusive sensors, is gaining increasing attention. Nevertheless, multi-modal 3D HPE still faces challenges, including modality imbalance. In this work, we introduce a novel balanced multi-modal learning method for 3D HPE, which harnesses the power of RGB, LiDAR, mmWave, and WiFi. Specifically, we propose a Shapley value-based contribution algorithm to assess the contribution of each modality and detect modality imbalance. To address this imbalance, we design a modality learning regulation strategy that decelerates the learning process during the early stages of training. We conduct extensive experiments on the widely adopted multi-modal dataset, MM-Fi, demonstrating the superiority of our approach in enhancing 3D pose estimation under complex conditions. We will release our codes soon.
@@ 1 Introduction
3D human pose estimation (3D HPE) recovers 3D coordinates of human joints from various input sources. It has gained significant research attention due to its applications in human-robot interaction [garcia2019human, zimmermann20183d], and computer animation [mehta2017vnect]. Specifically, in the rehabilitation context, where a variety of sensors and monitoring devices are deployed in the surroundings to detect patient action, 3D HPE models can act as an indispensable tool for supervising and verifying the correctness of patients’ exercises, ensuring adherence to established standards. Existing methods primarily focus on camera-based inputs (i.e., RGB images and videos) due to their accessibility and abundant human body information. However, camera-based approaches encounter limitations under occlusion scenarios and necessitate complex spatial conversion from 2D to 3D, dependent on accurate camera parameters. Hence multi-modal human sensing emerges as a promising approach for addressing complex scenarios by leveraging diverse sensor modalities. Wearable sensors are constrained by user compliance, hindering their practical adoption in everyday scenarios. Therefore, non-intrusive sensors such as LiDAR, mmWave radar, and WiFi offer advantages in terms of illumination invariance and user convenience, as shown in Figure 1 (a). By fusing information from RGB and non-intrusive sensors, we can enhance downstream task performance by exploiting both complementary and redundant information [wang2015mmss]. Nevertheless, existing methods [chen2023immfusion, an2022fast, he2024video, zheng2022multi, furst2021hperl] primarily rely on one or two modalities, leaving three or more modalities as an unexplored area.
@@
Figure 1: (a) Illustration of the multi-modal 3D HPE. (b) In end-to-end training, modality imbalance arises where dominant modalities with higher scores suppress the optimization of others in $M$-Fi [yang2024mm].
@@
Although multi-modal models are designed for extracting information from all modalities [qi2025robust, qi2025action, wang2024adversarial, deng2025global, lv2025t2sg, zhu2023unsupervised, qi2021semantics, qi2020stc, qi2019attentive], dominant modalities suppress the optimization of less dominant modalities in multi-modal learning as noted in [wang2020makes, peng2022balanced], leading to modality imbalance, shown in Figure 1 (b). While modality modulation techniques [peng2022balanced, li2023boosting] mitigate this issue by interfering with the learning of the dominant, they can degrade overall performance in certain scenarios, as discussed in [wei2024diagnosing]. This is attributed to their neglect of the intrinsic limitations of different modalities’ information capacity, resulting in a failure to achieve an effective balance. Consequently, the challenge lies in balancing modality optimizations without compromising the learning of dominant modalities.
@@
Furthermore, existing balancing methods [wu2022characterizing, peng2022balanced, li2023boosting, ReconBoost, fan2023pmr] are specifically tailored for discriminative tasks that rely on cross-entropy loss or involve explicit class attribution, making them inherently unsuitable for regression tasks. In addition, several approaches [wei2024innocent, guo2024classifier, wang2020makes] introduce auxiliary uni-modal components, such as uni-modal heads, to facilitate optimization. However, this design choice typically results in increased model complexity and additional optimization for those auxiliary parameters.
@@
First, we employ a Shapley value-based algorithm to quantify each modality’s contribution during training. Unlike previous balancing techniques that rely on auxiliary uni-modal heads or additional network branches, the Shapley analysis [shapley1953value] is fully parameter-free and can be directly computed within existing multi-modal fusion architectures. This enables the model to dynamically monitor the relative importance of modalities, serving as a reliable indicator of their optimization status and guiding adaptive balancing throughout training.
Second, we introduce a weight constraint loss that leverages the Fisher Information Matrix (FIM) [fisher1925theory] to adaptively regulate parameter updates based on their empirical importance. The FIM provides a principled, data-driven estimate of parameter sensitivity by approximating the expected squared gradient magnitude, which is an indicator of how much each parameter contributes to loss reduction. Crucially, in multimodal settings, dominant modalities tend to produce larger gradients in the beginning of training [huang2025adaptive, achille2018critical, huang2023fuller] and thus higher FIM values, while weaker modalities yield smaller ones. By penalizing deviations from initial parameters in proportion to their FIM-weighted importance, our loss applies stronger regularization to dominant modalities and milder constraints to inferior ones. This adaptive mechanism mitigates optimization imbalance: it curbs overly aggressive updates from strong modalities without excessively suppressing weaker ones, thereby fostering more balanced and robust multi-modal learning.
@@
(4) We conduct extensive experiments on the largest multi-modal 3D HPE dataset, MM-Fi [yang2024mm], demonstrating the superiority of our approach over baseline methods.
@@ 2 Related Work
3D Human Pose Estimation. Currently, the predominant focus of 3D HPE is on vision-based methods, categorized into two classes: directly estimating 3D joints [tekin2016direct, pavlakos2017coarse, wehrbein2021probabilistic, moon2019camera] and 2D-to-3D lifting [cai2019exploiting, peng2024ktpformer, li2024hourglass, xu2024finepose]. Recently, LiDAR has been applied to 3D HPE due to its robustness, which is utilized in autonomous driving [zheng2022multi, cong2022stcrowd]. MmWave-based HPE has gained increasing attention in research community [zhu2024probradarm3f], driven by the demand for privacy-preserving technologies. WiFi-based sensing has also emerged as a promising area of research [arshad2017wi, zou2018robust, geng2022densepose] due to its popularity. Although RGB-based methods are the majority in research community, multi-modal methods could sacrifice the computational complexity for better performance and robustness. In this work, we aim to integrate these modalities to enhance 3D HPE performance.
Multi-Modal Learning. Multi-modal learning has gained significant attention in recent years. Several works [wang2020makes, peng2022balanced, guo2024classifier] focus on end-to-end multi-modal training, revealing issues like modality imbalance or competition, causing the performance inferior to that of uni-modal models [peng2022balanced]. While another methods [li2023boosting, peng2022balanced] are proposed to quantify modality imbalance and improve training by modulating the gradients of dominant modalities. PMR [fan2023pmr] introduces a prototype learning approach based on class-specific representations in classification tasks, followed by modality-specific optimization strategies. MMPareto [wei2024innocent] incorporates the concept of the Pareto front into multi-modal learning and optimizes the multi-modal gradient based on joint uni-modal gradients. Although existing methods have made significant progress in mitigating modality imbalance, most are tailored to classification tasks and often rely on additional learnable parameters. In contrast, our work addresses modality imbalance in the context of regression, specifically 3D human pose estimation, by leveraging Shapley value-based contribution assessment and an adaptive regularization term to enable balanced optimization. Crucially, our approach introduces no extra learnable parameters.
@@ 3.2 Shapley Value-Based Contribution Algorithm
Shapley Value [shapley1953value] was introduced in coalition game theory to address profit distribution by calculating each player’s marginal contribution to the group, ensuring fair distribution. Prior works on modality imbalance [hu2022shape, li2023boosting, wei2024enhancing] have introduced the Shapley value into multimodal learning to assess modality contributions in classification tasks but limited in the regression. Specifically, the contribution score $\phi^{m}$ for modality $m$ is obtained by calculating the profits provided by all permutations of $\mathcal{M}\setminus\{m\}$ cooperated with $m$. Let $S$ denotes a subset of $\mathcal{M}\setminus\{m\}$, the calculation is defined as:
@@
Shapley Value in classification tasks. In classification settings, the cross-entropy loss is commonly used as the profit function $s(\cdot,\cdot)$ in prior methods [hu2022shape, li2023boosting, wei2024enhancing]. However, in regression tasks, directly applying the MSE or MAE loss between ground truths $y$ and predictions $\hat{y}$ as the profit function becomes problematic.
@@
Figure 3: (a) Mean and (b) standard deviation of human joint coordinate predictions sampled from MM-Fi [yang2024mm] during the training process. Notably, inferior modalities, such as mmWave and WiFi, exhibit near-zero standard deviations, indicating limited variability in their predictions.
@@ 3.3 Adaptive Weight Constraint Regularization
Existing methods [wei2024innocent, huang2025adaptive] typically modulate either the direction or the magnitude of gradients in isolation, despite the fact that both jointly govern how modality-specific encoders are updated during training. Moreover, these approaches often focus exclusively on information-sufficient (i.e., dominant) modalities and overlook the risk of overfitting to noisy signals in weaker ones.
@@
Learning window. Furthermore, following prior work [huang2025adaptive, achille2018critical] showing that most task-relevant information is acquired during early training, we apply $\mathcal{L}_{\text{AWC}}$ only within this initial phase. In practice, the first $K$ epochs are defined as the learning window, which is a tunable hyperparameter. Restricting $\mathcal{L}_{\text{AWC}}$ to this window mitigates the rapid dominance of strong modalities in early optimization, giving weaker modalities sufficient opportunity to acquire useful representations and achieve balanced multimodal learning.
@@
+\ + G-Blending [wang2020makes] 58.40 37.20 53.48 34.30 53.13 33.28
+\ + OGM-GE [peng2022balanced] 55.51 35.92 52.37 32.58 51.68 32.84
+\ + AGM [li2023boosting] 55.80 38.10 54.21 37.95 53.88 36.30
+\ + Modality-level [wei2024enhancing] 53.24 34.81 55.31 30.88 53.98 31.85
@@
+\ + G-Blending [wang2020makes] 57.14 38.98 53.85 34.52 53.93 34.36
+\ + OGM-GE [peng2022balanced] - - - - - -
+\ + AGM [li2023boosting] 58.40 39.91 61.28 41.24 53.60 36.86
+\ + Modality-level [wei2024enhancing] 53.33 35.23 50.48 31.47 50.71 33.46
@@ 3.4 Training Processes and Objectives
The whole learning process of our method is outlined in Algorithm 1. The MPJPE loss function [pavllo20193d] $\mathcal{L}_{\text{MPJPE}}$ is utilized to optimize 3D HPE task:
@@ 4.1 Experimental Setup
Dataset. MM-Fi [yang2024mm] is the first multi-modal non-intrusive 4D human dataset, including four wireless sensing modalities: RGB-D, LiDAR, mmWave radar, and WiFi. This dataset comprises 1,080 video clips, totaling 320k synchronized frames, performed by 40 volunteers engaged in 14 daily activities and 13 rehabilitation exercises. We conduct the experiments on three scenarios outlined in [yang2024mm] based on activity categories, where Protocol 1 includes 14 daily activities, Protocol 2 includes 13 rehabilitation exercises, and Protocol 3 involves all activities. The S1 random split strategy in [yang2024mm] is adopted in our experiments.
Evaluation Metrics. In our experiments, we follow previous works [pavllo20193d] using the following evaluation protocols: Mean Per Joint Position Error (MPJPE) and Procrustes Analysis MPJPE (PA-MPJPE) in millimeters. And MPJPE is formulated as follows:
@@
Compared Methods. We compare our method with a joint-training baseline, MM-Fi [yang2024mm], X-Fi [chen2025xfi], HPE-Li [d2024hpe], and several modality balancing methods, including G-Blending [wang2020makes], OGM-GE [peng2022balanced], AGM [li2023boosting] and Modality-level resample [wei2024enhancing], across various fusion strategies: concatenation, MLP, and self-attention.
Implementation Details. We implement our methods based on Pytorch on two NVIDIA RTX 3090 GPUs. Following [yang2024mm], we utilize VideoPose3D [pavllo20193d] as the backbone for RGB modality, Point Transformer [zhao2021point] for LiDAR and mmWave, and MetaFi++ [zhou2023metafi++] for WiFi. For the multi-modal learning, we train the model from scratch with randomly initialized parameters in an end-to-end manner, without loading any pre-trained parameters. We employ the Adam optimizer with a momentum of 0.9, a weight decay of 1e-4 and a batch size of 192. The training epoch is 50. The initial learning rate is 1e-3, which is multiplied by 0.1 every 30 epochs.
@@ 5 Conclusion
[…]
Supplementary Material
6 Methodology Details
6.1 Network Architecture
Figure 7: Illustration of our model architecture.
As shown in Fig. 7, our model integrates RGB frames (R), LiDAR point clouds (L), mmWave point clouds (M), and WiFi CSI data (W) as multi-modal inputs for 3D human pose estimation. Modality-specific encoders are initially employed to extract features from the respective data sources. Specifically, VideoPose3D [pavllo20193d] is utilized as the backbone for RGB frames. 2D joint sequences of 81 frames are fed into the RGB encoder as input, where each frame comprises 17 joints with two coordinates (i.e., $x$ and $y$). The model dimension is set to 1024 following [pavllo20193d]. Point Transformer [zhao2021point] is leveraged for LiDAR and mmWave data, with the layer number $L$ set to 5 and the dimension fixed at 512. The sample stride is set to 1 for the first LiDAR layer and 4 for the subsequent four layers, whereas it remains constant at 1 for all mmWave layers. The input channels of the LiDAR and mmWave encoders are 3 and 5, respectively. MetaFi++ [zhou2023metafi++] is adopted for WiFi, with a dimension of 512. Subsequently, a multi-modal fusion module fuses features from these modality-specific branches, followed by a pose regression head (a linearlinear layer) to predict the final joint positions. Furthermore, four 2-layer MLPs are appended to the respective encoders, each containing a $LayerNorm$ and a $ReLU$ activation function, to unify feature dimensions across modalities. The output dimension of this 2-layer MLP is 256.
Three fusion strategies are employed in our experiments: concatenation, concatenation with MLP, and self-attention, to derive the fused feature $f_{fuse}$. For concatenation, features from different modalities are directly concatenated along the channel:
$f_{fuse}^{concat}=[f^{R}\ f^{L}\ f^{M}\ f^{W}]\in\mathbb{R}^{b\times(4\times d)},$ (12)
where $f^{m}$ represents the feature extracted by the encoder for modality $m$, with dimensions $b\times d$. Here, $b$ denotes the batch size and $d$ is the feature dimension, set to 256. In the concatenation with MLP strategy, a 3-layer MLP is applied to the concatenated features:
$f_{fuse}^{mlp}=mlp(Concat(f^{m}))\in\mathbb{R}^{b\times d},$ (13)
where $d$, the output dimension of the fusion MLP, is also set to 256. The first two MLP layers are followed by $ReLU$ activation functions. The self-attention strategy utilizes a 4-layer transformer encoder [vaswani2017attention] to fuse the features:
$s_{concat}$ $=[f^{R}\ f^{L}\ f^{M}\ f^{W}]\in\mathbb{R}^{b\times 4\times d}$ (14)
$s_{fuse}$ $=encoder(s_{concat})\in\mathbb{R}^{b\times 4\times d},$
$f_{fuse}^{attn}$ $=avgpool(s_{fuse})\in\mathbb{R}^{b\times d},$
where $s_{concat}$ is the sequence formed by concatenating features along the middle dimension. The final fused feature, $f_{fuse}$, is obtained by applying average pooling to the output of the transformer encoder.
Protocol 1 Protocol 2 Protocol 3
Methods MPJPE PA-MPJPE MPJPE PA-MPJPE MPJPE PA-MPJPE
Concat.+MLP 54.58 36.00 53.75 32.60 50.57 33.37
+\ + G-Blending [wang2020makes] 64.13 43.73 54.91 35.67 55.84 38.48
+\ + OGM-GE [peng2022balanced] - - - - - -
+\ + AGM [li2023boosting] 57.35 40.52 56.23 38.01 55.06 36.88
+\ + Modality-level [wei2024enhancing] 58.74 38.78 57.10 35.43 54.28 32.92
+\ + Ours 52.73 35.56 52.69 32.10 49.57 32.66
Table 5: Comparisons of our proposed method and existing balancing multi-modal learning methods on $M$-Fi under the fusion strategy of concatenation with MLP. The lower is better; the best results are highlighted in bold.
6.2 Shapley Value Computation
In this section, we present an illustrative example of the contribution score calculation process for the RGB modality:
$\phi^{R}(\mathcal{M})$ $=\frac{1}{4}[s(R,L,M,W)-s(0,L,M,W)$ (15)
$+s(R,0,0,0)]$
$+\frac{1}{12}[s(R,L,0,0)-s(0,L,0,0)$
$+s(R,0,M,0)-s(0,0,M,0)$
$+s(R,0,0,W)-s(0,0,0,W)]$
$+\frac{1}{12}[s(R,L,M,0)-s(0,L,M,0)$
$+s(R,L,0,W)-s(0,L,0,W)$
$+s(R,0,M,W)-s(0,0,M,W)],$
where $\mathcal{M}$ is the set of modalities and $s()$ is the Pearson correlation.
7 Experiments Details
Results of balanced multi-modal learning under concatenation with MLP. As shown in Table 5, our method effectively balances multi-modal optimization and achieves the best performance across all three protocols. In Protocol 1, our method obtains MPJPE and PA-MPJPE values of 52.73 mm and 35.56 mm, outperforming the baseline Concat.+MLP (54.58 mm and 36.00 mm) by 1.85 mm (3.39%) and 0.44 mm (1.22%), respectively. More notably, our approach consistently surpasses all competing methods, including G-Blending [wang2020makes], AGM [li2023boosting], and Modality-level [wei2024enhancing], in both metrics under all protocols. Specifically, in Protocol 2, our method achieves 52.69 mm (MPJPE) and 32.10 mm (PA-MPJPE), and in Protocol 3, it further improves to 49.57 mm and 32.66 mm, demonstrating superior performance.
Qualitative results. We present qualitative evaluations of our method in Figures 8 and 9. As shown in Figure 8, our approach achieves significantly better alignment with the ground truth compared to AGM [li2023boosting], particularly in dynamic postures involving limb articulation and body orientation. For example, in the first row, our method correctly reconstructs bent arms and rotated torsos, whereas AGM suffers from joint jittering and structural distortion. Moreover, Figure 9 illustrates how our model integrates diverse modalities, including RGB, LiDAR, mmWave, and WiFi, to produce robust and semantically consistent predictions.
Figure 8: Visual comparisons of 3D human pose estimation between AGM and our method on $M$-Fi.
Figure 9: Visualization of input modalities (RGB, LiDAR, mmWave, and WiFi) and their predictions.

### v5

@@ Abstract
3D human pose estimation (3D HPE) has emerged as a prominent research topic, particularly in the realm of RGB-based methods. However, the use of RGB images is often limited by issues such as occlusion and privacy constraints. Consequently, multi-modal sensing, which leverages non-intrusive sensors, is gaining increasing attention. Nevertheless, multi-modal 3D HPE still faces challenges, including modality imbalance. In this work, we introduce a novel balanced multi-modal learning method for 3D HPE, which harnesses the power of RGB, LiDAR, mmWave, and WiFi. Specifically, we propose a Shapley value-based contribution algorithm to assess the contribution of each modality and detect modality imbalance. To address this imbalance, we design a modality learning regulation strategy that decelerates the learning process during the early stages of training. We conduct extensive experiments on the widely adopted multi-modal dataset, MM-Fi, demonstrating the superiority of our approach in enhancing 3D pose estimation under complex conditions. Our source code is available at https://github.com/MICLAB-BUPT/AWC.
@@ 1 Introduction
3D human pose estimation (3D HPE) recovers 3D coordinates of human joints from various input sources. It has gained significant research attention due to its applications in human-robot interaction [13, 62], action assessment [43] and computer animation [26]. Specifically, in the rehabilitation context, where a variety of sensors and monitoring devices are deployed in the surroundings to detect patient action, 3D HPE models can act as an indispensable tool for supervising and verifying the correctness of patients’ exercises, ensuring adherence to established standards. Existing methods primarily focus on camera-based inputs (i.e., RGB images and videos) due to their accessibility and abundant human body information. However, camera-based approaches encounter limitations under occlusion scenarios and necessitate complex spatial conversion from 2D to 3D, dependent on accurate camera parameters. Hence multi-modal human sensing emerges as a promising approach for addressing complex scenarios by leveraging diverse sensor modalities. Wearable sensors are constrained by user compliance, hindering their practical adoption in everyday scenarios. Therefore, non-intrusive sensors such as LiDAR, mmWave radar, and WiFi offer advantages in terms of illumination invariance and user convenience, as shown in Figure 1 (a). By fusing information from RGB and non-intrusive sensors, we can enhance downstream task performance by exploiting both complementary and redundant information [47]. Nevertheless, existing methods [5, 2, 16, 59, 12] primarily rely on one or two modalities, leaving three or more modalities as an unexplored area.
@@
Figure 1: (a) Multi-modal 3D HPE. (b) In end-to-end training, modality imbalance arises where dominant modalities with higher scores suppress the training of others in $M$-Fi [56].
@@
Although multi-modal models are designed for extracting information from all modalities, dominant modalities suppress the optimization of less dominant modalities in multi-modal learning as noted in [49, 31], leading to modality imbalance, shown in Figure 1 (b). While modality modulation techniques [31, 21] mitigate this issue by interfering with the learning of the dominant, they can degrade overall performance in certain scenarios, as discussed in [53]. This is attributed to their neglect of the intrinsic limitations of different modalities’ information capacity, resulting in a failure to achieve an effective balance. Consequently, the challenge lies in balancing modality optimizations without compromising the learning of dominant modalities.
@@
Furthermore, existing balancing methods [54, 31, 21, 18, 10] are specifically tailored for discriminative tasks that rely on cross-entropy loss or involve explicit class attribution, making them inherently unsuitable for regression tasks. In addition, several approaches [52, 15, 49] introduce auxiliary uni-modal components, such as uni-modal heads, to facilitate optimization. However, this design choice typically results in increased model complexity and additional optimization for those auxiliary parameters.
@@
First, we employ a Shapley value-based algorithm to quantify each modality’s contribution during training. Unlike previous balancing techniques that rely on auxiliary uni-modal heads or additional network branches, the Shapley analysis [45] is fully parameter-free and can be directly computed within existing multi-modal fusion architectures. This enables the model to dynamically monitor the relative importance of modalities, serving as a reliable indicator of their optimization status and guiding adaptive balancing throughout training.
Second, we introduce a weight constraint loss that leverages the Fisher Information Matrix (FIM) [11] to adaptively regulate parameter updates based on their empirical importance. The FIM provides a principled, data-driven estimate of parameter sensitivity by approximating the expected squared gradient magnitude, which is an indicator of how much each parameter contributes to loss reduction. Crucially, in multimodal settings, dominant modalities tend to produce larger gradients in the beginning of training [19, 1, 20] and thus higher FIM values, while weaker modalities yield smaller ones. By penalizing deviations from initial parameters in proportion to their FIM-weighted importance, our loss applies stronger regularization to dominant modalities and milder constraints to inferior ones. This adaptive mechanism mitigates optimization imbalance: it curbs overly aggressive updates from strong modalities without excessively suppressing weaker ones, thereby fostering more balanced and robust multi-modal learning.
@@
(4) We conduct extensive experiments on the largest multi-modal 3D HPE dataset, MM-Fi [56], demonstrating the superiority of our approach over baseline methods.
@@ 2 Related Work
3D Human Pose Estimation. Video data plays a central role in many computer vision tasks [38, 36, 39, 41, 32, 42] and has become a widely used modality for 3D HPE. Existing approaches can be broadly categorized into two classes: directly estimating 3D joints [46, 28, 50, 27], and 2D-to-3D lifting [4, 30, 22, 55]. Recently, LiDAR has been applied to 3D HPE due to its robustness, which is utilized in autonomous driving [59, 7]. MmWave-based HPE has gained increasing attention in research community [61], driven by the demand for privacy-preserving technologies. WiFi-based sensing has also emerged as a promising area of research [3, 63, 14] due to its popularity. Although RGB-based methods are the majority in research community, multi-modal methods could sacrifice the computational complexity for better performance and robustness. In this work, we aim to integrate these modalities to enhance 3D HPE performance.
Multi-Modal Learning. Multi-modal learning [44, 33, 23, 37, 40] has attracted significant attention across diverse perception and understanding tasks [24, 25, 9, 48, 57, 34, 35]. However, training such models remains challenging due to complex interactions among modalities. Several works [49, 31, 15] focus on end-to-end training, revealing issues like modality imbalance or competition, causing the suboptimal performance [31]. To address this issue, several methods [21, 31] are proposed to quantify modality imbalance and improve training by modulating the gradients of dominant modalities. PMR [10] introduces a prototype learning approach based on class-specific representations in classification tasks. MMPareto [52] incorporates the concept of the Pareto front into multi-modal learning and optimizes the multi-modal gradient based on joint uni-modal gradients. Although existing methods have made significant progress in mitigating such imbalance, most are tailored to classification tasks and often rely on additional learnable parameters. In contrast, our work addresses modality imbalance in the context of regression, specifically 3D human pose estimation, by leveraging Shapley value-based contribution assessment and an adaptive regularization to enable balanced optimization. Crucially, our approach introduces no extra learnable parameters.
@@ 3.2 Shapley Value-Based Contribution Algorithm
Shapley Value [45] was introduced in coalition game theory to address profit distribution by calculating each player’s marginal contribution to the group, ensuring fair distribution. Prior works on modality imbalance [17, 21, 51] have introduced the Shapley value into multimodal learning to assess modality contributions in classification tasks but limited in the regression. Specifically, the contribution score $\phi^{m}$ for modality $m$ is obtained by calculating the profits provided by all permutations of $\mathcal{M}\setminus\{m\}$ cooperated with $m$. Let $S$ denotes a subset of $\mathcal{M}\setminus\{m\}$, the calculation is defined as:
@@
Shapley Value in classification tasks. In classification settings, the cross-entropy loss is commonly used as the profit function $s(\cdot,\cdot)$ in prior methods [17, 21, 51]. However, in regression tasks, directly applying the MSE or MAE loss between ground truths $y$ and predictions $\hat{y}$ as the profit function becomes problematic.
@@
Figure 3: (a) Mean and (b) standard deviation of human joint coordinate predictions sampled from MM-Fi [56] during the training process. Modalities like mmWave and WiFi show near-zero standard deviations, indicating their predictions collapse to nearly constant values and contribute little to pose estimation.
@@ 3.3 Adaptive Weight Constraint Regularization
Existing methods [52, 19] typically modulate either the direction or the magnitude of gradients in isolation, despite the fact that both jointly govern how modality-specific encoders are updated during training. Moreover, these approaches often focus exclusively on information-sufficient (i.e., dominant) modalities and overlook the risk of overfitting to noisy signals in weaker ones.
@@
Learning window. Furthermore, following prior work [19, 1] showing that most task-relevant information is acquired during early training, we apply $\mathcal{L}_{\text{AWC}}$ only within this initial phase. In practice, the first $K$ epochs are defined as the learning window, which is a tunable hyperparameter. Restricting $\mathcal{L}_{\text{AWC}}$ to this window mitigates the rapid dominance of strong modalities in early optimization, giving weaker modalities sufficient opportunity to acquire useful representations and achieve balanced multimodal learning.
@@
+\ + G-Blending [49] 58.40 37.20 53.48 34.30 53.13 33.28
+\ + OGM-GE [31] 55.51 35.92 52.37 32.58 51.68 32.84
+\ + AGM [21] 55.80 38.10 54.21 37.95 53.88 36.30
+\ + Modality-level [51] 53.24 34.81 55.31 30.88 53.98 31.85
@@
+\ + G-Blending [49] 57.14 38.98 53.85 34.52 53.93 34.36
+\ + OGM-GE [31] - - - - - -
+\ + AGM [21] 58.40 39.91 61.28 41.24 53.60 36.86
+\ + Modality-level [51] 53.33 35.23 50.48 31.47 50.71 33.46
@@ 3.4 Training Processes and Objectives
The whole learning process of our method is outlined in Algorithm 1. The MPJPE loss function [29] $\mathcal{L}_{\text{MPJPE}}$ is utilized to optimize 3D HPE task:
@@ 4.1 Experimental Setup
Dataset. MM-Fi [56] is the first multi-modal non-intrusive 4D human dataset, including four wireless sensing modalities: RGB-D, LiDAR, mmWave radar, and WiFi. This dataset comprises 1,080 video clips, totaling 320k synchronized frames, performed by 40 volunteers engaged in 14 daily activities and 13 rehabilitation exercises. We conduct the experiments on three scenarios outlined in [56] based on activity categories, where Protocol 1 includes 14 daily activities, Protocol 2 includes 13 rehabilitation exercises, and Protocol 3 involves all activities. The S1 random split strategy in [56] is adopted in our experiments.
Evaluation Metrics. In our experiments, we follow previous works [29] using the following evaluation protocols: Mean Per Joint Position Error (MPJPE) and Procrustes Analysis MPJPE (PA-MPJPE) in millimeters. And MPJPE is formulated as follows:
@@
Compared Methods. We compare our method with a joint-training baseline, MM-Fi [56], X-Fi [6], HPE-Li [8], and several modality balancing methods, including G-Blending [49], OGM-GE [31], AGM [21] and Modality-level resample [51], across various fusion strategies: concatenation, MLP, and self-attention.
Implementation Details. We implement our methods based on Pytorch on two NVIDIA RTX 3090 GPUs. Following [56], we utilize VideoPose3D [29] as the backbone for RGB modality, Point Transformer [58] for LiDAR and mmWave, and MetaFi++ [60] for WiFi. For the multi-modal learning, we train the model from scratch with randomly initialized parameters in an end-to-end manner, without loading any pre-trained parameters. We employ the Adam optimizer with a momentum of 0.9, a weight decay of 1e-4 and a batch size of 192. The training epoch is 50. The initial learning rate is 1e-3, which is multiplied by 0.1 every 30 epochs.
@@ 5 Conclusion
6 Acknowledgement
This work is partly supported by the Funds for the National Natural Science Foundation of China under Grant 62572072, Beijing Natural Science Foundation under Grant L243027 and NSFC under Grant U24B20176.